# PPC-GPT: Federated Task-Specific Compression of Large Language Models via Pruning and Chain-of-Thought Distillation

### v1

@@ Abstract
Compressing Large Language Models (LLMs) into task-specific Small Language Models (SLMs) encounters two significant challenges: safeguarding domain-specific knowledge privacy and managing limited resources. To tackle these challenges, we propose PPC-GPT, a innovative privacy-preserving federated framework specifically designed for compressing LLMs into task-specific SLMs via pruning and Chain-of-Thought (COT) distillation. PPC-GPT works on a server-client federated architecture, where the client sends differentially private (DP) perturbed task-specific data to the server’s LLM. The LLM then generates synthetic data along with their corresponding rationales. This synthetic data is subsequently used for both LLM pruning and retraining processes. Additionally, we harness COT knowledge distillation, leveraging the synthetic data to further improve the retraining of structurally-pruned SLMs. Our experimental results demonstrate the effectiveness of PPC-GPT across various text generation tasks. By compressing LLMs into task-specific SLMs, PPC-GPT not only achieves competitive performance but also prioritizes data privacy protection.
@@
Tao Fan1, 2, Guoqiang Ma2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1The Hong Kong University of Science and Technology, Hong Kong, China 2WeBank Co., Ltd, Shenzhen, China Correspondence: tfanac@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 OpenAI (2023a) and LLaMA3-70B Touvron et al. (2023a), boasting billions of parameters and remarkable text generation capabilities, have emerged as a transformative force in the realm of artificial intelligence. However, their training demands substantial computational resources OpenAI (2023b), and their colossal size poses significant hurdles for practical deployment, especially in resource-limited environments. Conversely, Small Language Models (SLMs), such as OPT-1.3B Zhang et al. (2022) and Pythia-1.4B Biderman et al. (2023), frequently demonstrate superior computational efficiency and accelerated response rates, making them ideally suited for real-time applications with constrained resources. Enterprises with constrained resources typically prefer deploying SLMs, as they can do so without the concern of potential data leaks, a risk that is heightened when utilizing remote LLMs. Yet, training an SLM from scratch, even the smallest billion-parameter models, entails considerable computational expenses that are financially prohibitive for most enterprises. Furthermore, SLMs exhibit inherent limitations that stem from their performance constraints.
In this work, we aim to tackle the following question: Is it feasible to develop a task-specific and competitive SLM by harnessing an existing pre-trained LLM for enterprises with limited resources, while ensuring compliance with privacy requirements? To achieve this objective, we delve into structured pruning Xia et al. (2023); Men et al. (2024); Kim et al. (2024), as a viable approach. Pruning is generally regarded as a strategy for compressing task-specific models by eliminating redundant parameters and expediting inference, all while maintaining task performance.
@@
As depicted in Figure 1(a), the envisioned architecture of PPC-GPT comprises a high-performance server adept at deploying LLMs and facilitating their pruning into SLMs, coupled with a client endowed with more constrained computational capabilities for running SLMs. Within the confines of our framework, the workflow unfolds as detailed below. Initially, the client sends task-specific data, perturbed to ensure privacy, to the server. These data are protected by the Exponential Mechanism of Differential Privacy Dwork (2006); McSherry and Talwar (2007); Tong et al. (2023), thereby guaranteeing privacy protection. Subsequently, the server-side $LLM_{syn}$ generates synthetic data along with their corresponding rationales, based on these perturbed inputs. The server-side $LLM_{o}$, which represents the original model, undergoes pruning by PPC-GPT to yield the target $SLM_{t}$. This pruning process is informed by both the synthetic data and their associated rationales. Following the pruning of the $LLM_{o}$, the server retrains the target $SLM_{t}$ through COT  Wei et al. (2022); Hsieh et al. (2023); Li et al. (2023) knowledge distillation, leveraging the same synthetic data and rationales. Lastly, the server dispatches the refined target $SLM_{t}$ to the client, who then proceeds to retrain the $SLM_{t}$ utilizing its locally private data.
@@
  • Privacy-Preserving Federated LLM Compression Framework. We propose PPC-GPT, an innovative privacy-preserving federated framework tailored for compressing LLMs into task-specific SLMs. We use LLM to create synthetic data from DP-perturbed task data. This data helps evaluate layer importance in LLM, guiding the pruning process.
  • COT Knowledge Distillation Leveraging Synthetic Data. We incorporate COT knowledge distillation, utilizing synthetic data generated from the LLM. This approach enhances the retraining of structurally-pruned SLMs, thereby improving their overall performance.
  • Empirical Assessment of LLM Compressing to Task-Specific SLM. Through comprehensive compressing experiments conducted on the LLaMA and OPT models across a range of text generation tasks, PPC-GPT has demonstrated its efficacy in compressing LLMs into task-specific SLMs, achieving competitive performance.
(a) Overview of our proposed PPC-GPT framework.
(b) Layer Importance Example: The significance of each layer, as indicated by the BI (Block Influence) value of LLaMa2-7B on the OBQA dataset, based on the PPC-GPT framework.
Figure 1: The overview of our proposed PPC-GPT. The PPC-GPT comprises four key components: (1) The DP Perturbed Data, which perturbs the client’s data to ensure privacy; (2) The Synthetic Data Generation, responsible for creating new synthetic data and rationales based on the perturbed data; (3) The Layer-Wise Structured Pruning, a process that prunes original $LLM_{o}$ to obtain the target smaller $SLM_{t}$; (4) The Retraining SLM, where the target $SLM_{t}$ is retrained using both synthetic and original private data to restore accuracy.
@@ 2.1 Differential Privacy
$\epsilon$-Differential Privacy (DP). The Definition of $\epsilon$-Differential Privacy (DP) Dwork (2006). A randomized algorithm $M:D\to S$ is $\epsilon$-Differential Privacy if for any two neighboring datasets $D_{1},D_{2}\in D$ that differ exactly in a single data sample, and for any output $O\subseteq S$:
@@
Exponential Mechanism. The Definition of Exponential Mechanism McSherry and Talwar (2007); Tong et al. (2023). For a given scoring function $u:X\times Y\to R$, a randomized mechanism $M(X,u,Y)$ is $\epsilon$-DP compliant if it satisfies:
@@ 2.2 Differential Privacy Synthetic Data
A practical approach to generating private synthetic data involves training a language model, such as LLaMa2-7B Touvron et al. (2023a), on private data using DP through DP-SGD Song et al. (2013); Bassily et al. (2014); Abadi et al. (2016). Subsequently, the DP model is sampled repeatedly to produce synthetic data  Mattern et al. (2022); Yue et al. (2022); Kurakin et al. (2023). Research conducted by Mattern et al. (2022); Yue et al. (2022); Kurakin et al. (2023) demonstrates that training downstream models on DP synthetic data achieves performance comparable to training directly on real data with DP, thereby underscoring the high quality of the synthetic data.
However, a significant challenge arises because cutting-edge LLMs, like GPT-4, do not offer model weights, making DP fine-tuning impractical. Even for open-source LLMs, such as LLaMa3-70B Touvron et al. (2023a), the process is resource-intensive. Meanwhile, these DP fine-tuning methods inherently rely on a trusted server to gather data from data owners for model training Chen et al. (2023), significantly limiting their applicability in scenarios where such trusted servers are not available, as is the case in our research context. In the context of this work, we operate within a client-server architecture where fine-tuning the LLM on the server is not an option.
@@ 2.3 Model Pruning
Model pruning, initially proposed by LeCun et al. (1989) and subsequently enhanced by Han et al. (2015), stands as a resilient and efficient strategy for mitigating model redundancy and attaining compression. This methodology branches into two primary techniques: unstructured pruning and structured pruning.
Unstructured pruning  Dong et al. (2017); Lee et al. (2019); Wang et al. (2020); Sun et al. (2023); Frantar and Alistarh (2023) can obtain highly compressed models by directly pruning neurons, disregarding the model’s internal architecture, which also causes unstructured sparsity and hard deployment. A more pragmatic and structured option is structured pruning. Structured pruning targets organized patterns for removal, encompassing entire layers Jha et al. (2023), attention heads within Multi-Head Attention (MHA) mechanisms Michel et al. (2019), hidden sizes in Feedforward Neural Networks (FFN) Nova et al. (2023), as well as hybrid configurations Kurtić et al. (2024). In recent times, there has been a surge in structured pruning research tailored specifically for LLMs. For example, ShortGPT Men et al. (2024), LaCo Yang et al. (2024), and Shortened LLaMa Kim et al. (2024) concentrate solely on pruning depth (i.e., layer-wise). LLM-Pruner Ma et al. (2023) eliminates coupled structures in relation to network width while preserving the layer count. Sheared-LLaMA Xia et al. (2023) introduces a mask learning phase that is designed to pinpoint prunable components in both network width and depth. Our work falls in the category of structured pruning of LLMs.
@@ 3 The Proposed PPC-GPT Framework
In this section, we introduce PPC-GPT, an innovative privacy-preserving federated framework specifically designed for compressing LLMs into task-specific SLMs via pruning and COT distillation. We illustrate the PPC-GPT in Figure 1(a). The PPC-GPT comprises four key modules: DP Perturbed Data, Synthetic Data Generation, Layer-Wise Structured Pruning, and Retraining. We will elaborate on these modules in Section 3.2, Section 3.3, Section 3.4 and Section 3.5, respectively, following presenting the problem formulation we try to address in Section 3.1. Our aim is to provide a comprehensive understanding of how PPC-GPT addresses the intricate challenges associated with compressing LLMs while maintaining the client’s data privacy and optimizing task-specific performance.
3.1 Problem Formulation
Given an LLM $f_{\theta_{o}}$ with parameters $\theta_{o}$, which represents the original LLM that requires pruning, and a task-specific dataset $\mathcal{D}$ containing private data, our objective is to develop a target smaller, task-specific compressed SLM $f_{\theta_{t}}$ parameterized by $\theta_{t}$. To acheive this, we seek to find the optimal pruning strategy $\mathcal{P}$ and retraining approach $\mathcal{R}$. The objective can be formulated as follows:
$\quad\quad\quad\quad\min_{\mathcal{P},\mathcal{R}}\mathcal{L}(\theta_{t};\theta_{o},\mathcal{D})$ (4)
$s.t.\quad\left|\theta_{t}\right|\ll\left|\theta_{o}\right|\ \quad\text{and}\quad\mathcal{L}_{p}(\mathcal{D})<\delta$
where $\mathcal{L}(\theta_{t};\theta_{o},\mathcal{D})$ is the loss function measuring the performance of the compressed SLM on the task-specific dataset. $|\theta_{t}|$ and $|\theta_{o}|$ denote the number of parameters in the compressed and original models, respectively. $\mathcal{L}_{p}(\mathcal{D})$ is the privacy loss incurred due to the perturbation of the data to ensure differential privacy.
3.2 DP Perturbed Data
We utilize an exponential mechanism McSherry and Talwar (2007); Yue et al. (2021); Chen et al. (2023) to perturb the local private data $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{N}$, which satisfies the criteria for the $\epsilon$-DP. For detailed information about the exponential mechanism, please refer to Section 2.1. We denote the perturbed dataset as $\mathcal{D}_{p}=\left\{(x^{p}_{i})\right\}_{i=1}^{N}$, where $x^{p}_{i}$ signifies an perturbed input based on the original local private dataset $\mathcal{D}$
The Exponential Mechanism $\mathcal{M}$ is defined as a randomized algorithm that, given the original local private dataset $\mathcal{D}$, outputs the perturbed dataset $\mathcal{D}_{p}$ with probability proportional to the exponential of the utility score:
@@ 3.2 DP Perturbed Data
3.3 Synthetic Data Generation
When the server-side $LLM_{syn}$ receives the perturbed data $\mathcal{D}_{p}$, the server initiates a procedure where $LLM_{syn}$ generates fresh synthetic data along with their corresponding rationales based on these perturbed data. We denote the synthetic dataset as $\mathcal{D}_{s}=\left\{(x_{i},(y_{i},r_{i}))\right\}_{i=1}^{N_{s}}$, where $x_{i}$ signifies an input, $y_{i}$ signifies the corresponding expected output label, $r_{i}$ signifies the desired rationale, and $N_{s}$ represents the sample size of synthetic data.
We introduce a simple and efficient method for generating synthetic data, utilizing prompt engineering techniques and CoT technology:

### v2

@@ Abstract
Compressing Large Language Models (LLMs) into task-specific Small Language Models (SLMs) encounters two significant challenges: safeguarding domain-specific knowledge privacy and managing limited resources. To tackle these challenges, we propose PPC-GPT, a novel unified framework that systematically addresses both privacy preservation and model compression in federated settings. PPC-GPT works on a server-client federated architecture, where the client sends differentially private (DP) perturbed task-specific data to the server’s LLM. The LLM then generates synthetic data along with their corresponding rationales. This synthetic data is subsequently used for both LLM pruning and retraining processes. Our framework’s key innovation lies in its holistic integration of privacy-preserving mechanisms, synthetic data generation, and task-specific compression techniques, creating unique benefits through component interaction. Our experiments across diverse text generation tasks demonstrate that PPC-GPT successfully achieves dual objectives: maintaining competitive performance comparable to full-sized LLMs while ensuring robust privacy protection through its federated architecture. Our code has been contributed to the FATE open-source project and is now publicly accessible at https://github.com/FederatedAI/FATE-LLM/tree/main/python/fate_llm/algo/ppc-gpt
@@
Tao Fan1, 2, Guoqiang Ma2, Yuanfeng Song2, Lixin Fan2, Qiang Yang3 1Hong Kong University of Science and Technology, Hong Kong, China 2WeBank Co., Ltd, Shenzhen, China 3Hong Kong Polytechnic University, Hong Kong, China Correspondence: tfanac@cse.ust.hk, qyang@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 Gpt-4 and LLaMA3-70B dubey2024llama, boasting billions of parameters and remarkable text generation capabilities, have emerged as a transformative force in the realm of artificial intelligence. However, their training demands substantial computational resources openai2023gpt4tr, and their colossal size poses significant hurdles for practical deployment, especially in resource-limited environments. Conversely, Small Language Models (SLMs), such as OPT-1.3B zhang2022opt and Qwen2.5-1.5B team2024qwen2, frequently demonstrate superior computational efficiency and accelerated response rates, making them ideally suited for real-time applications with constrained resources. Enterprises with constrained resources typically prefer deploying SLMs, as they can do so without the concern of potential data leaks, a risk that is heightened when utilizing remote LLMs fan2025ten; fan2025fedmkt; fan2023fate-llm; kang2023grounding. Yet, training an SLM from scratch, even the smallest billion-parameter models, entails considerable computational expenses that are financially prohibitive for most enterprises. Furthermore, SLMs exhibit inherent limitations that stem from their performance constraints.
In this work, we aim to tackle the following question: Is it feasible to develop a task-specific and competitive SLM by harnessing an existing pre-trained LLM for enterprises with limited resources, while ensuring compliance with privacy requirements? To achieve this objective, we delve into structured pruning xia2024sheared; men2024shortgpt; kim2024shortened, as a viable approach. Pruning is generally regarded as a strategy for compressing task-specific models by eliminating redundant parameters and expediting inference, all while maintaining task performance.
@@
As depicted in Figure 1, the envisioned architecture of PPC-GPT comprises a high-performance server adept at deploying LLMs and facilitating their pruning into SLMs, coupled with a client endowed with more constrained computational capabilities for running SLMs. Within the confines of our framework, the workflow unfolds as detailed below. Initially, the client sends task-specific data, perturbed to ensure privacy, to the server. These data are protected by the Exponential Mechanism of Differential Privacy dwork2006differential; mcsherry2007mechanism; tong2025inferdpt, thereby guaranteeing privacy protection. Subsequently, the server-side auxiliary $LLM_{syn}$ generates synthetic data along with their corresponding rationales, based on these perturbed inputs. The server-side $LLM_{o}$, which represents the original model, undergoes pruning by PPC-GPT to yield the target SLM. This pruning process is informed by both the synthetic data and their associated rationales. Following the pruning of the $LLM_{o}$, the server retrains the target SLM through CoT  wei2022chain; hsieh2023distilling; li2023symbolic knowledge distillation, leveraging the same synthetic data and rationales. Lastly, the server dispatches the refined target SLM to the client, who then proceeds to retrain the target SLM utilizing its locally private data.
@@
  • Unified Framework for Federated LLM Compression. We propose PPC-GPT, the first holistic framework designed to systematically unify privacy-preserving techniques with LLM compression in a federated learning context. Its core innovation lies in the synergistic integration of four key modules: exponential mechanism-based data perturbation, CoT-guided synthetic data generation, rationale-aware structured pruning, and CoT-based knowledge distillation. This modular architecture is not only novel but also extensible, allowing for future advancements in any of its constituent parts.
  • Novel Rationale-Aware Structured Pruning for Reasoning Preservation. We propose a new structured pruning metric that evaluates network layers based on their contribution to generating explanatory rationales. This approach selectively preserves the model’s essential reasoning capabilities, a critical aspect often overlooked in standard compression techniques.
  • Component Interaction Benefits. The integration creates unique advantages through component interaction. For example, combining DP-perturbed data with CoT-guided synthetic data generation enables both privacy protection and effective knowledge transfer, while the rationale-aware structured pruning leverages this enhanced data quality for better compression decisions.
  • Empirical Assessment of LLM Compressing to Task-Specific SLM. Through extensive experiments across various text generation tasks using LLaMA and OPT models, we demonstrate how PPC-GPT’s component interactions lead to effective task-specific compression while maintaining privacy, achieving results competitive with full-sized LLMs.
Figure 1: The overview of our proposed PPC-GPT. The PPC-GPT comprises four key components: (1) The Exponential Mechanism-based Data Perturbation, which perturbs the client’s data to ensure privacy; (2) The CoT-Guided Synthetic Data Generation, responsible for creating new synthetic data and rationales based on the perturbed data; (3) The Rationale-Aware Structured Pruning, a process that prunes original $LLM_{o}$ to obtain the target smaller SLMSLM; (4) The Retraining SLM, where the target SLMSLM is retrained using both synthetic and original private data to restore accuracy.
@@ 2.1 Differential Privacy
$\epsilon$-Differential Privacy (DP). The Definition of $\epsilon$-Differential Privacy (DP) dwork2006differential. A randomized algorithm $M:D\to S$ is $\epsilon$-Differential Privacy if for any two neighboring datasets $D_{1},D_{2}\in D$ that differ exactly in a single data sample, and for any output $O\subseteq S$:
@@
Exponential Mechanism. The Definition of Exponential Mechanism mcsherry2007mechanism; tong2025inferdpt. For a given scoring function $u:X\times Y\to R$, a randomized mechanism $M(X,u,Y)$ is $\epsilon$-DP compliant if it satisfies:
@@ 2.2 Differential Privacy Synthetic Data
A practical approach to generating private synthetic data involves training a language model, such as LLaMA2-7B touvron2023llama2, on private data using DP through DP-SGD song2013stochastic; bassily2014private; abadi2016deep. Subsequently, the DP model is sampled repeatedly to produce synthetic data  mattern2022differentially; yue-etal-2023-synthetic; kurakin2023harnessing. Research conducted by mattern2022differentially; yue-etal-2023-synthetic; kurakin2023harnessing demonstrates that training downstream models on DP synthetic data achieves performance comparable to training directly on real data with DP, thereby underscoring the high quality of the synthetic data.
However, a significant challenge arises because cutting-edge LLMs, like GPT-4, do not offer model weights, making DP fine-tuning impractical. Even for open-source LLMs, such as LLaMA3-70B dubey2024llama, the process is resource-intensive. Meanwhile, these DP fine-tuning methods inherently rely on a trusted server to gather data from data owners for model training chen-etal-2023-customized, significantly limiting their applicability in scenarios where such trusted servers are not available, as is the case in our research context. In the context of this work, we operate within a client-server architecture where fine-tuning the LLM on the server is not an option.
@@ 2.3 Model Pruning
Model pruning, initially proposed by lecun1989optimal and subsequently enhanced by han2015learning, stands as a resilient and efficient strategy for mitigating model redundancy and attaining compression. This methodology branches into two primary techniques: unstructured pruning and structured pruning.
Unstructured pruning  dong2017learning; lee2018snip; wang2020picking; sun2024a; frantar2023sparsegpt can obtain highly compressed models by directly pruning neurons, disregarding the model’s internal architecture, which also causes unstructured sparsity and hard deployment. A more pragmatic and structured option is structured pruning. Structured pruning targets organized patterns for removal, encompassing entire layers jha2023train, attention heads within Multi-Head Attention (MHA) mechanisms michel2019sixteen, hidden sizes in Feedforward Neural Networks (FFN) nova2023gradient, as well as hybrid configurations kurtic2024ziplm. In recent times, there has been a surge in structured pruning research tailored specifically for LLMs. For example, ShortGPT men2024shortgpt, LaCo yang2024laco, and Shortened LLaMA kim2024shortened concentrate solely on pruning depth (i.e., layer-wise). LLM-Pruner ma2023llm eliminates coupled structures in relation to network width while preserving the layer count. Sheared-LLaMA xia2024sheared introduces a mask learning phase that is designed to pinpoint prunable components in both network width and depth. Our work falls in the category of structured pruning of LLMs.
3 Problem Formulation
Given an LLM $f_{\theta}$ with parameters $\theta$, which represents the original LLM that requires pruning, and a task-specific dataset $\mathcal{D}$ containing private data, our objective is to develop a target smaller, task-specific compressed SLM $f_{\phi}$ parameterized by $\phi$. To acheive this, we seek to find the optimal pruning strategy $\mathcal{P}$ and retraining approach $\mathcal{R}$. The objective can be formulated as follows:
$\quad\quad\quad\quad\min_{\mathcal{P},\mathcal{R}}\mathcal{L}(\phi;\theta,\mathcal{D})$ (4)
$s.t.\quad\left|\phi\right|\ll\left|\theta\right|\ \quad\text{and}\quad\mathcal{L}_{p}(\mathcal{D})<\delta$
where $\mathcal{L}(\phi;\theta,\mathcal{D})$ is the loss function measuring the performance of the compressed SLM on the task-specific dataset. $|\phi|$ and $|\theta|$ denote the number of parameters in the compressed and original models, respectively. $\mathcal{L}_{p}(\mathcal{D})$ is the privacy loss incurred due to the perturbation of the data to ensure differential privacy.
@@ 4 The Proposed PPC-GPT Framework
In this section, we introduce PPC-GPT, a unified privacy-preserving federated framework for compressing LLMs into task-specific SLMs. We illustrate the PPC-GPT architecture in Figure 1. As detailed in Algorithm 1, our framework comprises four key modules that work in concert:
  • Exponential Mechanism-based Data Perturbation: Ensures privacy protection through exponential mechanism.
  • CoT-guided Synthetic Data Generation: Creates high-quality synthetic data and rationales.
  • Rationale-Aware Structured Pruning: Leverages synthetic data and rationales for model compression.
  • Retraining SLM: Optimizes the compressed model through two-stage knowledge distillation.
We elaborate on these modules in Section 4.1, 4.2, 4.3 and 4.4, respectively. Through this integrated approach, PPC-GPT effectively addresses the challenges of privacy-preserving LLM compression while maintaining task-specific performance.
Algorithm 1 PPC-GPT
0: Private dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, $LLM_{syn}$ for synthetic data generation, Original $LLM_{o}$ $f_{\theta}$ that requires pruning, Privacy budget $\epsilon$
0: Task-specific SLM $f_{\phi}$ with privacy guarantees
1:Phase 1: Exponential Mechanism-based Data Perturbation
2: Apply Exponential Mechanism $\mathcal{M}$ to $\mathcal{D}$ with budget $\epsilon$
3: Generate $\mathcal{D}_{p}=\{(x^{p}_{i})\}_{i=1}^{N}$ according to Eq.(5)
4:Phase 2: CoT-guided Synthetic Data Generation
5: Server’s $LLM_{syn}$ processes $\mathcal{D}_{p}$ to generate synthetic data
6: Generate $\mathcal{D}_{s}=\{(x^{s}_{i},(y^{s}_{i},r^{s}_{i}))\}_{i=1}^{N_{s}}$
7:Phase 3: Rationale-Aware Structured Pruning
8: Calculate Block Influence scores using $\mathcal{D}_{s}$ according to Eq.(7)
9: Identify redundant layers based on BI values
10: Obtain pruned model structure $f_{\phi}$ where $|\phi|\ll|\theta|$
11:Phase 4: Retraining SLM via Two-Stage Knowledge Distillation
12: Server performs CoT distillation using $\mathcal{D}_{s}$ according to Eq.(8), (9), (10)
13: Client fine-tunes with private data $\mathcal{D}$ according to Eq.(11)
14:return Compressed task-specific SLM $f_{\phi}$
4.1 Exponential Mechanism-based Data Perturbation
We utilize an exponential mechanism mcsherry2007mechanism; yue2021differential; chen-etal-2023-customized to perturb the local private data $\mathcal{D}=\left\{(x_{i},y_{i})\right\}_{i=1}^{N}$, which satisfies the criteria for the $\epsilon$-DP. For detailed information about the exponential mechanism, please refer to Section 2.1. We denote the perturbed dataset as $\mathcal{D}_{p}=\left\{(x^{p}_{i})\right\}_{i=1}^{N}$, where $x^{p}_{i}$ signifies an perturbed input based on the original local private dataset $\mathcal{D}$ .
The Exponential Mechanism $\mathcal{M}$ is defined as a randomized algorithm that, given the original local private dataset $\mathcal{D}$, outputs the perturbed dataset $\mathcal{D}_{p}$ with probability proportional to the exponential of the utility score (in this work, we use cosine similarity as the utility function):
@@ 4.1 Exponential Mechanism-based Data Perturbation
4.2 CoT-guided Synthetic Data Generation
When the server-side $LLM_{syn}$ receives the perturbed data $\mathcal{D}_{p}$, the server initiates a procedure where $LLM_{syn}$ generates fresh synthetic data along with their corresponding rationales based on these perturbed data. We denote the synthetic dataset as $\mathcal{D}_{s}=\left\{(x^{s}_{i},(y^{s}_{i},r^{s}_{i}))\right\}_{i=1}^{N_{s}}$, where $x^{s}_{i}$ signifies an input, $y^{s}_{i}$ signifies the corresponding expected output label, $r^{s}_{i}$ signifies the desired rationale, and $N_{s}$ represents the sample size of synthetic data.
As illustrated in Figure 2, we introduce a simple and efficient method for generating synthetic data, utilizing prompt engineering techniques and CoT technology: